OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9

OpenStudy (shane_b):

I'd use the substitution method I just showed you on your other question.

OpenStudy (anonymous):

can you show me please

OpenStudy (anonymous):

\[x-2y=14\] \[x+3y=9\] substract top from bottom get 5y=-5 divide by 5 y=-1 so what is x?

OpenStudy (anonymous):

how would you get x?

OpenStudy (shane_b):

Look at @kmalone99's post...he solved it for y already. So you know what y is...just plug it into one of the original equations and solve for x: x − 2y = 14 x - 2(-1)=14

OpenStudy (anonymous):

so 12?

OpenStudy (anonymous):

yes very good

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

great job